OpenStudy (shelld93):

Can someone walk me through 3 1/5 divided by 8/9

OpenStudy (luigi0210):

Convert the fractions: \[\LARGE 3\frac{1}{5}=\frac{16}{5}\] And.. \[\LARGE \frac{a}{b} \div \frac{c}{d}=\frac{a}{b}*\frac{d}{c}\] So.. \[\LARGE \frac{16}{5}*\frac{9}{8}\]

OpenStudy (shelld93):

how did you get 16 ?

OpenStudy (luigi0210):

Now cross cancel.. 16/8=2: \[\LARGE \frac{2}{5}*\frac{9}{1}\]